OpenStudy (anonymous):

I'm having trouble following the limit of x/sin 3x as x approaches 0 explanation

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{x}{\sin(3x)}=\lim_{x \rightarrow 0}\frac{3}{3} \cdot \frac{x}{\sin(3x)}=\lim_{x \rightarrow 0}\frac{1}{3} \cdot \frac{3x}{\sin(3x)}=\frac{1}{3} \cdot 1\]

myininaya (myininaya):

satellite do you need me to do 1/3 * 1?

OpenStudy (anonymous):

please i have trouble with fractions

myininaya (myininaya):

lol

myininaya (myininaya):

\[\frac{1}{3} \cdot 1=\frac{1}{3}\]

OpenStudy (anonymous):

merci

OpenStudy (anonymous):

thanks!!